OpenStudy (anonymous):

solve: a^2 · a^3 · a^-4

OpenStudy (anonymous):

do you know the answer?

OpenStudy (tamtoan):

use these formula ..a^-n = 1/a^n , a^n . a^m = a^(n+m) and a^n/a^m = a^(n-m) and you can get the answer quick

OpenStudy (anonymous):

it's not a^9 though..

OpenStudy (tamtoan):

ofcourse it's not, one of the power is negative...check that again :)

OpenStudy (anonymous):

see that's what confuses me, but I thought when you flip the exponent to make it positive, you add it as a positive exponent?

OpenStudy (anonymous):

would it be a^1?

OpenStudy (tamtoan):

correct

OpenStudy (anonymous):

okay, So you're not supposed to flip it then? :S

OpenStudy (tamtoan):

it's two positive powers 2 and 3..you add..one negative power 4 ..you subtract

OpenStudy (anonymous):

ty

OpenStudy (tamtoan):

yw :)